Fermi National Laboratory

Accelerator Update for July 28, 2008 - July 30, 2008


Monday July 28

The day shift began with Operations monitoring the Tevatron (TeV) store 6320 (68.82E30) and the Recycler (RR) stash (320.1E10), with the Antiproton Source stacking (31.8E10), with the Main Injector, Booster, and Linac operating normally, with the Collider Detector at Fermilab (CDF) and DZero (D0) using the colliding beams, and with the SciBooNE / MiniBooNE (BNB), NuMI, and Meson MTest experiments taking beam.

Operators put Pbar into standby at 9:15 AM.

Operations halted the MI ramp at 9:18 AM to allow power supply experts to work on the MI-10 power supply.

Operations terminated store 6320 at 9:20 AM.

At 9:27 AM, a Booster RF station (BRF10) tripped off. It reset.

The MI power supply worked ended and Pbar resumed stacking at 9:36 AM.

Operations reported at 9:39 AM that they had prepared CDF and D0 for collision hall accesses.

At 9:50 AM, cryo system experts began replacing the TeV sector D1 wet engine connecting rod bearings.

An instrumentation expert began upgrading the TeV BPMs at 9:57 AM.

Operations disabled the MTest beam around 12:17 PM.

A MI expert began a slipstacking study at 1:38 PM.

At 2:43 PM, MIRF3 tripped off. An expert investigated and reset the station.

CDF came out of access at 2:52 PM.

The MI study ended at 3 PM.

Operations put the TeV through a dry squeeze at 3:53 PM.

Operations reported at 4 PM that the Recycler had stashed antiprotons three times during the day shift.

Operations put the TeV through a wet squeeze at 5:05 PM.

Operations put the TeV into shot setup at 5:31 PM.

Operations established store 6323 at 7:01 PM with an initial luminosity of 305.7E30.

At 7:52 PM, LRF4 tripped off. It reset.

Operations established beam to MTest at 8:04 PM.

At 10:36 PM, Operations reported that the MI-30 LCW temperatures were rising. Operations contacted an expert. An operator reduced the MI repetition rate to reduce temperatures.

Operations reported at midnight that the Recycler had stashed three times during the evening shift.

Tuesday July 29

The midnight shift began with Operations monitoring store 6323 (143.67E30) and the stash (205.0E10), with Pbar stacking (37.3E10), with the BNB, NuMI, and MTest experiments taking beam.

At 12:41 AM, Operations reported that a Linac Klystron RF station (KRF3) had tripped several times on a LCW flow fault indication. The station reset.

Operations increased the MI beam repetition rate around 12:45 AM.

At 6:41 AM, a MI vertical trim magnet power supply (I:V701) and a Lambertson (LAM52) tripped off and caused the beam permits for MI and the P1 transfer line to trip. Everything reset.

At 6:56 AM, LRF4 tripped off. It reset.

Operations reported at 8:07 AM that the Recycler had stashed five times during the midnight shift.

At 9:29 AM, Operations reported that MIRF11 had tripped off several times on a Bias Supply Shunt Regulator Overcurrent indication. This caused the NuMI beam permit to trip. Everything reset.

A NuMI expert reported at 9:46 AM that there was a planned power outage for the MINOS Sudan mine in Minnesota that would require NuMI to shut down for about two hours this afternoon.

Operations terminated store 6323 at 9:54 AM.

Operations disabled the MTest beam at 10:02 AM.

Operations put the TeV into shot setup at 10:04 AM.

Operations established store 6324 at 11:56 AM with an initial luminosity of 293.47E39.

Operators prepared the Pelletron for an access at 12:15 PM.

At 1:01 PM, KRF7 tripped off for the first of two times in an hour. It reset.

A Pbar expert began a Pbar Debuncher cooling measurement study at 1:05 PM.

At 1:14 PM, LRF 4 tripped off. Operators and experts responded. They had the station back on within 30 minutes.

Operations established beam to MTest around 2:30 PM.

Operations reported at 2:39 PM that there had been earthquakes in Papua New Guinea and Los Angeles. No TeV beam was lost.

(TeV experts discovered that the Tevatron could feel earthquakes when, after a TeV quench, they noticed that there was also an earthquake in Alaska at the same time. The above picture shows how the TeV tilt meters reacted to the latest earthquake.)

Operations disabled the NuMI beam at 3:11 PM for the Sudan mine power outage.

Operators prepared Pbar for access at 3:32 PM.

A MI expert began a MI-8 transfer line study at 3:54 PM.

Operations reported at 4:01 PM that the Recycler had stashed four times during the day shift.

The Pbar access ended at 5:21 PM.

At 5:45 PM, Operations reported that they could not turn on a Pbar Debuncher power supply (D:H717).

At 6:03 PM, Operators reported a water flow alarm for D:H717 and sensors verified an LCW leak.

At 6:33 PM, Operations accessed Pbar and reported a massive water leak on IB-7 and obvious signs of arcing. Operations contacted experts.

A Pbar power supply expert reported at 7:19 PM that the D:H717 power supply never ground faulted.

Booster experts accessed their tunnel and at 9:20 PM reported that they had repaired BRF2.

At 11:12 PM, Pbar experts reported three problems with IB-7, the last of a string of seven magnets that make up D:H717: the flag un-welded itself from the copper going into the magnet; there was a water leak on one of the upper nuts for the LCW connection; and there was some ground current on the cables between the magnet and the power supply.

The water experts reported at 11:26 PM that they could not fix the leaking nut. They will investigate it further in the morning. Pbar came out of access.

Wednesday July 30

The midnight shift began with Operations monitoring store 6324 (83.39E30), and the stash (92.9E10), with Pbar off, and with BNB, NuMI, and MTest experiments off.

Operations established beam to NuMI and BNB at 12:05 AM.

Operations established beam to MTest at 12:08 AM.

At 5:19 AM, Operations reported a flurry of TeV RF alarms all related to IRM node TRF10 in TeV sector F0. The alarms lasted for a couple of minutes and then cleared. Operations will inform the RF and controls experts in the morning.

The Plans for Wednesday

The plans for today are to try to fix the Pbar magnet H717 in place. The repair won’t be completed until this evening and then the magnet will have to be checked to make sure the repair didn’t do any damage. (Replacing the magnet would take about seven days.) The store will stay in as long as it remains useful.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/30/2008    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory